OpenStudy (anonymous):

At 37 °C the concentration of Fe3+ inside a cell is 0.053 M and outside is 0.044 M. The cell membrane is permeable to Fe3+. What potential difference in volts would have to exist across the membrane for Fe3+ to be in equilibrium at the stated conditions? Give you answer as the absolute value of the potential difference in volts.

OpenStudy (aaronq):

Use the nernst equation : \(E_{cell}=E-\dfrac{RT}{nF}ln\dfrac{[Fe^{3+}_{out}]}{[Fe^{3+}_{in}]}\) At eq. \(\Delta G = 0=-nFE_{cell} \rightarrow E_{cell} =0 \) and thus: \(E =\dfrac{RT}{nF}ln\dfrac{[Fe^{3+}_{out}]}{[Fe^{3+}_{in}]}\)